OpenStudy (anonymous):

this is soo frusterating.. the computer doesnt let me get past this assignment. help please. Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 9-inch perimeter A = _______ sq. in.

OpenStudy (anonymous):

\[s^2\frac{\text{Sqrt}[3]}{4}\]

OpenStudy (anonymous):

This will find you area of equilateral triangle, remember each side is 3

OpenStudy (anonymous):

so then i would end up with 2 and 1/4?

OpenStudy (anonymous):

\[\frac{9 \sqrt{3}}{4}\]

OpenStudy (anonymous):

Your s = 3, so 3 * 3 * Sqrt(3) all over 4 is your area.

OpenStudy (anonymous):

P=9 inches than each side = 3 inches A=(b*h)/2 3/2=1.5 h= 1.5tan60 h=2.5980 A=(3*2.6)/2 A=3.897114317 you round

OpenStudy (anonymous):

Don't worry. She will be back soon, saying, Computer won't accept my answer.

OpenStudy (anonymous):

Most homework website demand exact answer

OpenStudy (anonymous):

Actually, many demand, rounding off to a certain decimal point.

OpenStudy (anonymous):

thanks everyone . but the computer is lame and tells me im wrong. and that its incorrectD: chaguanas is right computers want exact answers :(

OpenStudy (anonymous):

Check out instructions, it might say round off to so many digits or something like that.